OpenStudy (anonymous):

An equation is shown below: 9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points) please help me i will fan you

OpenStudy (anonymous):

please help i really to finish this i will give you an invisible cupcake with a cherry on top :D

OpenStudy (anonymous):

okay thanks for the cooky well for part a i think you have to plug in random values of either x or y and find y or x resultants values and make 4 points from 4 plugged values

OpenStudy (anonymous):

actually thats for part 2

OpenStudy (anonymous):

okay ill do that then what about part b and c

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

i guess it works for both

OpenStudy (anonymous):

i think for a lets say x=c then y=9c+2 So all the possible solutions are (c,9c+2) (you can also keep x)

OpenStudy (anonymous):

what part is that for

OpenStudy (anonymous):

a

OpenStudy (anonymous):

okay thanks so much you get all the cupcakes in the world. what about part c

OpenStudy (anonymous):

well to find inconsistent solution just find an equation that is parallel to this one..meaning they have the same slope so[ y=9x+c (but c doesnt equal to +2 )] would be inconsistent solution

OpenStudy (anonymous):

oh i get it thank you sooooooooo much. do you think you could help me with one more

OpenStudy (anonymous):

yw and for sure dude

OpenStudy (anonymous):

k here it is A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

first of all you cookies taste wonderful :)

OpenStudy (anonymous):

aww thanks

OpenStudy (anonymous):

okay you can use elimination method. since -3x + 7y = -16 -9x + 5y = 16 \[\frac{ 9 }{ 3 }*\frac{-3x + 7y = -16 }{ -9x + 5y = 16 }\]

OpenStudy (anonymous):

and by simplify it and finding the value of y you can plug in one of the equations to find x

OpenStudy (anonymous):

yay thanks so much *gives a batch of chocolate cupcakes* enjoy!! :D

OpenStudy (anonymous):

ohhhhhhhhhhhhhh wait im sorry that is different solution i didnt read the last section of part one

OpenStudy (anonymous):

im sorry but this one is THE ONE look since we have the following equations -3x + 7y = -16 ---- (1) -9x + 5y = 16 ---- (2) and it says we have to "Replace one equation with the sum of that equation and a multiple of the other." Let us replace (2) by adding (2) to two times (1): 2 times (1): -6x + 14y = -32 Add (2): -9x + 5y = 16 -------------------- -15x + 19y = -16 So the second pair of equations are: -3x + 7y = -16 ---- (1) -15x + 19y = -16 ----- (3) For B) Solve (1) and (2) for x,y Solve (1) and (3) for x,y Show they are the same solutions. its very simple did you got it ?

OpenStudy (anonymous):

yes thanks i finally understand. the problem with online math classes is that they dont explain anything but you explained everything perfectly thanks so much

OpenStudy (anonymous):

yw